# The weak Extension Principle

### v1

@@ 2. Proof of the main result
$V_{n}^{i}$ =$=\bigcup\{\mathrm{supp}(\psi_{n}^{i}(h))\mid h\in C_{0}(U_{n}^{i})\}$
=$=\{x\in X\mid\exists h\in C_{0}(U_{n}^{i}),0\leq h\leq 1,\psi_{n}^{i}(h)(x)=1\}$
=$=\{x\in X\mid\exists h\in C_{0}(U_{n}^{i}),0\leq h\leq 1,\Psi(h)(x)=1\}.$
Hence, for each $k$ we can find positive contractions $h_{k}\in C_{0}(U_{n_{k}}^{i})$ such that
$\psi_{n_{k}}^{i}(h_{k})(x_{k})=\Psi(h_{k})(x_{k})=1.$
Since $x_{k}\notin\bigcup_{n\in S}V_{n}^{e}=\Gamma(\bigcup_{n\in S}U_{n}^{e})$, we can further assume that $\mathrm{supp}(h_{k})\cap\bigcup_{n\in S}U_{n}^{e}=\emptyset$.
Let $h=\sum h_{k}\in\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i})$, and notice that for every $T\subseteq{\mathbb{N}}$ we have $\Psi(h_{T})(x_{k})=1$ whenever $k$ is such that $n_{k}\in T$ and $h_{T}f_{S}=0$, since $\mathrm{supp}(f_{S})\subseteq\bigcup_{n\in S}U_{n}^{e}$ which is disjoint from $\mathrm{supp}(h)\supseteq\mathrm{supp}(h_{T})$.
@@
Since $F$ is a continuous map between compact spaces, it is closed, hence so is $F[U_{0}]$. Suppose for a contradiction that there is a nonempty open $O\subseteq F[U_{0}]\subseteq Y^{*}$. Let $O\subseteq\beta Y$ be open such that $O\cap Y^{*}=U$, and let $O^{\prime}=O\cap Y$. Since the sets $\{U_{n}^{i}\mid i\in\{e,o\},n\in{\mathbb{N}}\}$ cover $Y$, we can find an infinite $S\subseteq{\mathbb{N}}$ and $i\in\{e,o\}$ such that $O^{\prime}\cap U_{n}^{i}\neq\emptyset$ for all $n\in S$. For $n\in S$, let $h_{n}$ be a positive contraction whose support is in $U_{n}^{e}$ and such that there is $x\in O^{\prime}$ with $h_{n}(x)=1$. Let $h=\sum h_{n}$. Note that for every infinite $T\subseteq S$ we have that $\left\lVert\Phi(h_{T})\chi_{U_{0}}\right\rVert=1$. Find $T\subseteq S$ infinite such that $\Psi$ lifts $\Phi$ on $h_{T}$, and notice that $\chi_{U_{0}}\pi_{X}(\Psi(h_{T}))=0$. This is a contradiction. ∎
Claims 2.4 and 2.6 together show that $G$, $U_{0}$ and $U_{1}$ satisfy the hypotheses of $\operatorname{{\mathsf{wEP}}}(X,Y)$. This concludes the proof of the theorem. ∎
@@
Assume $\operatorname{{\mathsf{OCA}}}$ and $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$. Let $\kappa<\lambda$ be infinite cardinals, and suppose that $X$ and $Y$ are locally compact noncompact second countable spaces. Then there is no continuous surjection $(X^{*})^{\kappa}\to(Y^{*})^{\lambda}$.
@@
This is a consequence of Theorem 5.1 in [6], which asserts that the thesis holds if we have $\operatorname{{\mathsf{wEP}}}$. The result then follows from Theorem 1.2. ∎
The $\operatorname{{\mathsf{wEP}}}$ has consequences on the existence of surjections from Čech–Stone remainders to powers of them: while $\operatorname{\mathsf{CH}}$ implies that $\mathbb{N}^{*}$ surjects onto every compact space of density $\aleph_{1}$, and therefore onto every power of remainder of every (reasonable small) space, the same cannot happen if $\operatorname{{\mathsf{wEP}}}$ holds, where for example that $(\mathbb{N}^{*})^{d}$ surjects onto $(\mathbb{N}^{*})^{\ell}$ if and only if $d\geq\ell$ (see [5, §4.6]). In future work we intend to explore these phaenomena, along with extension principles in the noncommutative setting.

### v2

@@ 2. Proof of the main result
$V_{n}^{i}$ $=\bigcup\{\mathrm{supp}(\psi_{n}^{i}(h))\mid h\in C_{0}(U_{n}^{i})\}$
$=\{x\in X\mid\exists h\in C_{0}(U_{n}^{i}),0\leq h\leq 1,\psi_{n}^{i}(h)(x)=1\}$
$=\{x\in X\mid\exists h\in C_{0}(U_{n}^{i}),0\leq h\leq 1,\Psi(h)(x)=1\}.$
Hence, for each $k$ we can find positive contractions $h_{n_{k}}\in C_{0}(U_{n_{k}}^{i})$ such that
$\psi_{n_{k}}^{i}(h_{n_{k}})(x_{k})=\Psi(h_{k})(x_{k})=1.$
Since $x_{k}\notin\bigcup_{n\in S}V_{n}^{e}=\Gamma(\bigcup_{n\in S}U_{n}^{e})$, we can further assume that $\mathrm{supp}(h_{n_{k}})\cap\bigcup_{n\in S}U_{n}^{e}=\emptyset$.
Set $h_{j}=0$ if $j\notin\{n_{k}\mid k\in\mathbb{N}\}$, and let
$h=\sum h_{k}\in\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i}).$
Notice that for every $T\subseteq{\mathbb{N}}$ we have $\Psi(h_{T})(x_{k})=1$ whenever $k$ is such that $n_{k}\in T$ and $h_{T}f_{S}=0$, since $\mathrm{supp}(f_{S})\subseteq\bigcup_{n\in S}U_{n}^{e}$ which is disjoint from $\mathrm{supp}(h)\supseteq\mathrm{supp}(h_{T})$.
@@
Since $F$ is a continuous map between compact spaces, it is closed, hence so is $F[U_{0}]$. Suppose for a contradiction that there is a nonempty open $U\subseteq F[U_{0}]\subseteq Y^{*}$. Let $O\subseteq\beta Y$ be open such that $O\cap Y^{*}=U$, and let $O^{\prime}=O\cap Y$. Since the sets $\{U_{n}^{i}\mid i\in\{e,o\},n\in{\mathbb{N}}\}$ cover $Y$, we can find an infinite $S\subseteq{\mathbb{N}}$ and $i\in\{e,o\}$ such that $O^{\prime}\cap U_{n}^{i}\neq\emptyset$ for all $n\in S$. For $n\in S$, let $h_{n}$ be a positive contraction whose support is in $U_{n}^{e}$ and such that there is $x\in O^{\prime}$ with $h_{n}(x)=1$. Let $h=\sum h_{n}$. Note that for every infinite $T\subseteq S$ we have that $\left\lVert\Phi(h_{T})\chi_{U_{0}}\right\rVert=1$. Find $T\subseteq S$ infinite such that $\Psi$ lifts $\Phi$ on $h_{T}$, and notice that $\chi_{U_{0}}\pi_{X}(\Psi(h_{T}))=0$. This is a contradiction. ∎
Claims 2.5 and 2.7 together show that $G$, $U_{0}$ and $U_{1}$ satisfy the hypotheses of $\operatorname{{\mathsf{wEP}}}(X,Y)$. This concludes the proof of the theorem. ∎
@@
Assume $\operatorname{{\mathsf{OCA}}}$ and $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$. Let $\kappa<\lambda$ be cardinals, and suppose that $X$ and $Y$ are locally compact noncompact second countable spaces. Then there is no continuous surjection $(X^{*})^{\kappa}\to(Y^{*})^{\lambda}$.
@@
This is a consequence of Theorem 5.1 in [Fa:Dimension], which asserts that the thesis holds if we have $\operatorname{{\mathsf{wEP}}}$. The result then follows from Theorem 1.2. ∎
The $\operatorname{{\mathsf{wEP}}}$ has consequences on the existence of surjections from Čech–Stone remainders to powers of them: while $\operatorname{\mathsf{CH}}$ implies that $\mathbb{N}^{*}$ surjects onto every compact space of density $\aleph_{1}$, and therefore onto every power of remainder of every (reasonable small) space, the same cannot happen if $\operatorname{{\mathsf{wEP}}}$ holds, where for example that $(\mathbb{N}^{*})^{d}$ surjects onto $(\mathbb{N}^{*})^{\ell}$ if and only if $d\geq\ell$ (see [Fa:AQ, §4.6]). In future work we intend to explore these phenomena, along with extension principles in the noncommutative setting.